# From Fuzzy to Exact: The Halo Architecture for Infinite-Depth Reasoning via Rational Arithmetic

### v2

@@ Abstract
The pursuit of scale in deep learning has entrenched a trade-off: computational throughput is prioritized at the expense of numerical precision. We argue this compromise is fundamentally at odds with the requirements of general intelligence. We propose the Exactness Hypothesis: high-order causal reasoning—a cornerstone of AGI—demands a substrate supporting arbitrary-precision, logically consistent arithmetic.
We trace prevalent LLM failures, such as logical hallucinations and incoherence, to the inherent limitations of IEEE 754 floating-point arithmetic, where approximation errors compound catastrophically in deep functions. As a solution, we present the Halo Architecture, which transitions the computational foundation from approximate reals ($\mathbb{R}$) to exact rationals ($\mathbb{Q}$). Halo is realized through a custom Exact Inference Unit (EIU), whose design—featuring asynchronous MIMD reduction and dual-modular redundancy—resolves the performance and reliability bottlenecks of exact computation at scale.
@@ 1 Introduction
In this work, we propose the Halo Architecture, which moves from the approximate field of real numbers ($\mathbb{R}$) to the exact field of rational numbers ($\mathbb{Q}$). We ensure that operations such as $1/3+1/3+1/3$ yield exactly $1$, rather than $0.9999999$.
@@ 2 Related Work
Recent explorations into scaling "test-time compute," such as the Huginn-0125 prototype[Ren and Collaborators, 2025], have proposed decoupling parameter count from reasoning depth via latent loops. However, these attempts revealed a critical barrier: models relying on standard BF16 arithmetic suffer from "state collapse" (divergence) after approximately 50 recurrent steps. We view these findings not as architectural failures, but as empirical evidence of the "Precision Wall." Halo provides the necessary numerical substrate (Exact Rational Arithmetic) to stabilize these infinite-depth architectures, enabling them to fulfill their theoretical promise
@@ 3.2 The Derivative of Truth
3.3 Theoretical Proof: The Non-Associativity of Floats
@@
Consider a distributed training setup reducing gradients from 3 GPUs ($g_{1},g_{2},g_{3}$). If the reduction order changes (e.g., due to network latency), the resulting gradient update $\Delta\theta$ changes:
@@ 4.1 Algorithmic Formulation
The Halo architecture unifies infinite-precision arithmetic with recurrent semantic grounding. Algorithm 1 details the end-to-end inference process, strictly distinguishing between the “Light” stream (Rational $\mathbb{Q}$) and the “Ring” mechanism (Float $\mathbb{R}$).
Algorithm 1 The Halo Architecture: End-to-End Recurrent Inference
@@
1:Stage 1: Prelude (Embedding & Rationalization)
@@
4:Stage 2: The Light (Recurrent Core in EIU)
@@
6:  {Exact Arithmetic Transformer Block}
7:  $H_{attn}\leftarrow\text{RationalAttention}(H_{t-1})$
8:  $H_{mlp}\leftarrow\text{RationalFFN}(H_{attn})$
9:  $H_{temp}\leftarrow H_{mlp}+H_{t-1}$ {Lossless Residual Connection}
10:  Stage 3: The Ring (Conditional Activation)
11:  if $t\mod K==0$ then
12:   {Vector-Symbolic Collapse & Entropy Reset}
13:   $H_{t}\leftarrow\text{TheRing}(H_{temp})$
14:  else
15:   $H_{t}\leftarrow H_{temp}$
16:  end if
17:end for
18:Stage 4: Coda (Projection & Prediction)
@@
21:return $\text{Softmax}(logits)$
Algorithm 2 The Ring Mechanism & Rational Ops
@@
7: {3. Re-grounding to Rational Space (GCD Reduction)}
@@
16:return $A\cdot V$
17:End Function
4.2 The Light: Infinite Precision Stream
“The Light” represents the unattenuated flow of information. We define the computational medium as a stream of exact fractions $p/q$. Unlike floating-point numbers which drift, “The Light” ensures that the inference depth is limited only by time, not by precision loss.
4.3 The Ring: Vector-Symbolic Collapsing
To manage the potential complexity growth of rational numbers, we introduce The Ring. At every $N$-th cycle in the recurrent loop, the continuous vector $v$ is forced to collapse into discrete symbols (Language), verified for logical consistency, and then re-embedded.
@@ 4.3 The Ring: Vector-Symbolic Collapsing
This process resets the semantic entropy to zero, acting as a logical anchor.
The Ring mechanism can be viewed through the lens of the Information Bottleneck principle [Tishby et al., 2000]: by collapsing the state to a finite set of rationals, we explicitly filter out the “nuisance variables” (numerical noise) while retaining the “relevant variables” (logical structure).
4.4 The Ring as Bayesian Refinement
From a probabilistic perspective, the accumulation of floating-point noise represents a drift in the Likelihood $P(Observation|State)$. The Ring mechanism functions as a recursive Bayesian update step.
We define a discrete Rational Prior $\mathcal{P}(\theta)$ over the grid of simplest fractions $\mathbb{Q}_{grid}$. At every $K$-th step, the continuous, noisy state vector $v_{noisy}$ (the Observation) is combined with the Prior to compute the Maximum A Posteriori (MAP) estimate:
$v_{clean}=\arg\max_{q\in\mathbb{Q}_{grid}}P(v_{noisy}|q)P(q)$ (8)
The Ring effectively collapses the diffuse probability mass caused by numerical diffusion back into a Dirac delta function centered on the most likely rational state. This prevents the "uncertainty explosion" characteristic of chaotic dynamical systems.
4.5 Implementation: The Exact Inference Unit
Current GPUs are optimized for "Fuzzy Computing" with complex floating-point logic. The Halo Architecture relies on a new class of hardware, the Exact Inference Unit (EIU), which simplifies the datapath by reverting to fundamental integer arithmetic:
  1. 1.
    Integer-Only Datapath: The EIU eliminates complex Floating-Point Units (FPUs) entirely. Instead, it relies exclusively on standard integer multipliers, adders, and shifters to process rational numbers as pairs $(n,d)$. This simplifies the silicon footprint by removing costly exponent alignment and normalization circuits found in FP16/BF16 units.
  2. 2.
    Dynamic Bit-width: Registers that expand dynamically to accommodate the necessary precision for zero-loss calculation.
  3. 3.
    Deterministic Parallelism: Guaranteeing that $(A+B)+C=A+(B+C)$. This associativity is essential for consistent logical reasoning and is natively supported by the integer ring structure.
  4. 4.
    Hardware-Enforced Integrity (SDC Prevention): To support ultra-scale clusters ($10^{5}$ devices), the EIU incorporates Dual-Modular Redundancy directly into the datapath. This mechanism eliminates Silent Data Corruption (SDC) by verifying operations modulo Mersenne primes, ensuring exascale reliability (details in Appendix C).
  5. 5.
    Asynchronous MIMD Reduction: To ensure training throughput parity to current LLM, the EIU decouples simplification from arithmetic. Dedicated GCD engines (using Stein’s algorithm) operate asynchronously, masking the latency of reduction behind the main matrix multiplication pipeline (details in Appendix E).
4.6 Transcending the Irrational: Rational Approximations
Rational Softmax via Taylor Expansion. The standard attention mechanism requires computing $A=\text{softmax}(QK^{T}/\sqrt{d})$. We first replace the scaling factor $1/\sqrt{d}$ with an integer shift $2^{-k}$. For the exponential function, we utilize the property that for any $x\in\mathbb{Q}$, the truncated Taylor series yields a rational number:
$\text{RatExp}(x,N)=\sum_{k=0}^{N}\frac{x^{k}}{k!}\in\mathbb{Q}$ (9)
where $N$ is a dynamic precision parameter determined by the EIU. Unlike floating-point approximations which have fixed error, RatExp allows us to trade compute time for arbitrary precision. The Rational Softmax is then defined as:
Since both numerator and denominator are in $\mathbb{Q}$, the result remains strictly rational.
$y_{t+1}=\frac{1}{2}y_{t}\left(3-\sigma^{2}y_{t}^{2}\right)$ (11)
@@ 4.7 Training Objective: The Rational Alignment Loss
Unlike standard Vector Quantization (VQ) which maps vectors to an arbitrary learned codebook Van Den Oord et al. [2017], Halo maps vectors to the fixed algebraic field of logical rationals $\mathbb{Q}_{K}=\{\frac{n}{d}\mid|d|\leq K\}$.
@@
By minimizing $\mathcal{L}_{ring}$, the model learns to organize its latent space into a “grid” of precise rational numbers, effectively reducing the Information Bottleneck Tishby et al. [2000] and making the periodic “Ring Reset” operation lossless.
Our Rational Alignment Loss draws inspiration from the codebook commitment loss in VQ-VAE [Van Den Oord et al., 2017], but replaces the learned dictionary with a fixed, axiomatically true set of rational numbers. By minimizing $\mathcal{L}_{ring}$, the model learns to organize its latent space into a “grid” of precise rational numbers, effectively reducing the Information Bottleneck.
4.8 Formal Bound on Complexity Growth
A primary concern with Rational Arithmetic is the explosion of bit-width during repeated multiplication. We prove that the Halo architecture bounds this growth.
Theorem 1 (The Halo Boundedness Theorem). Let $B_{0}$ be the bit-width of the initial embedding. Let $K$ be the Ring Reset Interval. Let $\alpha$ be the maximum bit-growth rate per layer (for matrix multiplication, $\alpha\approx 2$). For any inference depth $L\to\infty$, the maximum bit-width $B_{max}$ required by the EIU is strictly bounded by:
$B_{max}\leq B_{ring}+K\cdot\alpha$ (13)
where $B_{ring}$ is the fixed bit-width of the Ring’s codebook.
Proof Sketch. In the “Light” phase, rational operations $(n_{1}/d_{1})\cdot(n_{2}/d_{2})$ result in a new denominator $d^{\prime}=d_{1}d_{2}$. In the worst case, bit-width grows linearly with depth $t$: $Bits(t)\approx Bits(t-1)+\alpha$. However, at step $t=mK$, the Ring mechanism projects the state $H_{t}$ onto the nearest rational in the fixed set $\mathbb{Q}_{K}$. This operation explicitly truncates the representation, resetting the bit-width to $B_{ring}$ (e.g., 64 bits). Thus, the maximum complexity occurs immediately before a reset, at step t=mK1t=mK-1. The accumulated growth is confined to the interval $[0,K]$. Consequently, the memory requirement is $\mathcal{O}(K)$, independent of total depth $L$. This transforms the complexity class of exact inference from $\mathcal{O}(e^{L})$ to $\mathcal{O}(1)$ amortized over infinite time. ∎
@@ 5.4 Scale Analysis: The 600B Parameter Regime
Results in Figure 5 reveal the “Curse of Dimensionality”: larger models are less numerically stable. This creates a paradox: **The larger the model, the higher the precision required.**
@@ 3. The Purity of Optimization
6.3 The Cost of Truth
Critics may argue that infinite precision is computationally intractable. We counter that while biological neural networks are noisy, the logical constructs humans create (mathematics, code) are exact. An AGI capable of generating novel physics or code must operate in a substrate that mimics this exactness. We are trading speed for truth.

### v3

@@ Abstract
The pursuit of scale in deep learning has entrenched a trade-off: computational throughput is prioritized at the expense of numerical precision. We argue this compromise is fundamentally at odds with the requirements of general intelligence. We propose the Exactness Hypothesis: high-order causal reasoning—a cornerstone of AGI—demands a substrate supporting arbitrary-precision, logically consistent arithmetic. We trace prevalent LLM failures, such as logical hallucinations and incoherence, to the inherent limitations of IEEE 754 floating-point arithmetic, where approximation errors compound catastrophically in deep functions.
As a solution, we present the Halo Architecture, which transitions the computational foundation from approximate reals ($\mathbb{R}$) to exact rationals ($\mathbb{Q}$). Halo is realized through a custom Exact Inference Unit (EIU), whose design—featuring asynchronous MIMD reduction and dual-modular redundancy—resolves the performance and reliability bottlenecks of exact computation at scale. Crucially, we theoretically prove that Halo’s number-theoretic quantization yields a quadratic error decay ($\mathcal{O}(D^{-2})$), strictly superior to the linear barrier of standard fixed-point arithmetic.
@@ 1 Introduction
In this work, we propose the Halo Architecture, which transitions the computational foundation from the approximate field of real numbers ($\mathbb{R}$) to the exact field of rational numbers ($\mathbb{Q}$). We ensure that operations such as $1/3+1/3+1/3$ yield exactly $1$, rather than $0.9999999$.
However, Halo goes beyond mere exactness. We introduce the Ring Mechanism to solve the tractability of rational arithmetic. Unlike standard quantization which truncates values linearly ($\mathcal{O}(D^{-1})$), our approach utilizes Diophantine Approximation to project states onto the Farey Sequence. We theoretically prove that this achieves a quadratic error decay ($\mathcal{O}(D^{-2})$), allowing Halo to preserve significantly higher information density per bit. We term this process Semantic Crystallization: using the discrete nature of number theory to filter out numerical noise while locking in the logical signal.
@@ 2 Related Work
Recent explorations into scaling "test-time compute," such as the Huginn-0125 prototype[Geiping et al., 2025], have proposed decoupling parameter count from reasoning depth via latent loops. However, these attempts revealed a critical barrier: models relying on standard BF16 arithmetic suffer from "state collapse" (divergence) after approximately 50 recurrent steps. We view these findings not as architectural failures, but as empirical evidence of the "Precision Wall." Halo provides the necessary numerical substrate (Exact Rational Arithmetic) to stabilize these infinite-depth architectures, enabling them to fulfill their theoretical promise
@@ 3.2 The Derivative of Truth
3.3 Algebraic Consistency: Determinism via Associativity
@@
If the reduction order changes (e.g., due to network latency), the resulting gradient update $\Delta\theta$ changes:
@@
3.4 Topological Density: Escaping the Sparsity-Error Loop
Critiques of low-precision formats often treat rounding error ($\epsilon$) and representational sparsity as distinct issues. We argue they are mutually exacerbating. Commodity floating-point formats (including IEEE 754 FP32/FP64 and the widely adopted BF16) can only represent a finite, non-uniform subset of the real number line $S_{FP}\subset\mathbb{R}$. The "Sparsity" of this grid directly dictates the magnitude of the rounding error.
It is a common misconception that migrating to Double Precision (FP64) solves this issue. While FP64 increases the density of the grid, it remains a sieve with structural holes. We quantify the Representational Capacity ($|S|$) of these formats compared to the Halo EIU (assuming 1024-bit dynamic registers):
$\frac{|S_{Halo}|}{|S_{FP64}|}\approx\frac{2^{2048}}{2^{64}}=2^{1984}\approx 10^{597}$ (7)
This staggering magnitude difference ($10^{597}$) implies that even FP64 covers an infinitesimally small fraction of the logical states expressible by Halo.
Crucially, the architecture of modern transformers ($x_{l+1}=x_{l}+F(x_{l})$) actively drives the state into these sparse regions. In high-dimensional geometry, residual updates tend to increase the vector norm $x||$. The deep learning community has empirically observed this magnitude explosion, identifying it as "Outlier Dimensions" [Dettmers et al., 2022] or the emergence of "Attention Sinks" [Xiao et al., 2023]. However, prior work typically frames these as semantic features or quantization nuisances. We reinterpret them as a symptom of the Sparsity-Error Trap: As $x||$ grows, the IEEE 754 exponent increases, causing the absolute gap between representable numbers to widen significantly (e.g., gap >0.01>0.01 near $10^{5}$). The model drifts into these regimes of lower precision, where the sparsity prevents the fine-grained subtraction required to correct the trajectory. Thus, error accumulation is not merely additive, but expansive—pushing the system irreversibly into the "fog" of low resolution.
@@ 4.1 Algorithmic Formulation
To address the precision-efficiency trade-off, Halo partitions the computational flow into two distinct regimes. We formally define these core components:
  • The Light (Exact Stream): The primary datapath where calculations occur in the arbitrary-precision rational field $\mathbb{Q}$ (represented as exact fractions $p/q$). In this stream, operations are chemically pure: no truncation or rounding occurs, ensuring $\epsilon=0$. Unlike floating-point numbers which drift, "The Light" ensures inference depth is limited only by time, not precision.
  • The Ring (Periodic Reset): A discrete bottleneck mechanism. It acts as a "Gatekeeper of Complexity," projecting the potentially complex rational state onto a fixed, low-bitwidth lattice to prevent the "explosion" of denominator complexity.
4.2 Halo-V: Vector-Symbolic Alignment (Baseline)
In our initial formulation (Halo-V), we implemented The Ring as a Vector-Symbolic Quantizer. This approach aligns the rational state with a statistically learned "Language of Thought" using VQ-VAE [Van Den Oord et al., 2017]. Algorithms 1 and 2 detail this baseline process, where the reset is triggered by a fixed schedule ($K$) and relies on learned Encoders/Decoders.
Algorithm 1 Halo-V: End-to-End Inference (Fixed Schedule)
@@ 4.2 Halo-V: Vector-Symbolic Alignment (Baseline)
1:Stage 1: Prelude
@@
4:Stage 2: The Light (Recurrent Core)
@@
6:  $H_{attn}\leftarrow\text{RationalAttention}(H_{t-1})$
7:  $H_{mlp}\leftarrow\text{RationalFFN}(H_{attn})$
8:  $H_{temp}\leftarrow H_{mlp}+H_{t-1}$
9:  Stage 3: The Ring (Fixed Schedule)
10:  if $t\mod K==0$ then
11:   {Vector-Symbolic Collapse}
12:   $H_{t}\leftarrow\text{TheRing}(H_{temp})$
13:  else
14:   $H_{t}\leftarrow H_{temp}$
15:  end if
16:end for
17:Stage 4: Coda
@@
20:return $\text{Softmax}(logits)$
Algorithm 2 Halo-V Ring Mechanism (Learned Quantizer)
@@
7: {3. Re-grounding to Rational Space}
@@
15:return $A\cdot V$
16:End Function
4.3 Implementation: The Exact Inference Unit (EIU)
To support the unique demands of Rational Arithmetic at scale, we introduce the Exact Inference Unit (EIU). Unlike GPUs optimized for approximate float math, the EIU features:
  1. 1.
    Integer-Only Datapath: The EIU eliminates FPUs entirely, relying exclusively on integer multipliers and adders to process rational pairs $(n,d)$.
  2. 2.
    Interrupt-Driven Registers: EIU registers are designed to operate in a free-running mode with dynamic bit-width. Crucially, they possess a Hardware Saturation Flag that triggers an interrupt strictly when the denominator’s MSB reaches the capacity limit (e.g., 4096 bits).
  3. 3.
    Asynchronous GCD Engines: Each Compute Unit is paired with a dedicated GCD engine capable of executing Stein’s Algorithm or Euclidean steps asynchronously from the main pipeline.
  4. 4.
    Dual-Modular Redundancy: To prevent Silent Data Corruption (SDC), operations are verified modulo Mersenne primes (see Appendix C).
4.4 Halo-N: Number-Theoretic Quantization via Diophantine Approximation
While Halo-V provides a robust logical baseline, the capabilities of the EIU enable a more advanced paradigm: Halo-N (Number-Theoretic). By leveraging the specific hardware features described above, Halo-N evolves the architecture from "Heuristic Approximation" to "Mathematical Optimality."
1. From Fixed Schedule to Natural Saturation
In Halo-V, the Ring is triggered at a fixed interval $K$. This is rigid. Halo-N utilizes the EIU’s Interrupt-Driven Registers to implement a "Natural Saturation" policy. The Ring reset is triggered only when the hardware flag asserts overflow. This transforms $K$ from a hyperparameter into a dynamic, physics-driven variable.
2. From VQ-VAE to Diophantine Solver
In Halo-V, The Ring relies on a learned Encoder/Decoder. Halo-N utilizes the EIU’s Asynchronous GCD Engines to replace this with a Diophantine Solver. Instead of learning a codebook, we project the state onto the Farey Sequence $\mathcal{F}_{N}$ using Coordinate-wise Diophantine Approximation. The mathematical process of finding the best rational approximation via Continued Fractions is algorithmically isomorphic to the Euclidean GCD algorithm, allowing the EIU to perform "Semantic Crystallization" without requiring any learned parameters.
Algorithms 3 and 4 detail this evolved process.
Algorithm 3 Halo-N: End-to-End Inference (Adaptive Saturation)
0: Input token sequence $X\in\mathbb{V}^{L}$
0: Recurrent Depth $T$ (Thinking Steps)
0: Saturation Limit $B_{max}$ (e.g., 4096 bits)
0: Next token probability distribution $P(y|X)$
1:Stage 1: Prelude
@@ 2. From VQ-VAE to Diophantine Solver
[…]
4:Stage 2: The Light (Recurrent Core)
5:for $t=1$ to $T$ do
6:  $H_{attn}\leftarrow\text{RationalAttention}(H_{t-1})$
7:  $H_{mlp}\leftarrow\text{RationalFFN}(H_{attn})$
8:  $H_{temp}\leftarrow H_{mlp}+H_{t-1}$
9:  Stage 3: The Ring (Interrupt-Driven Reset)
10:  if $\text{CheckSaturation}(H_{temp},B_{max})$ then
11:   {Triggered by Hardware Overflow Flag}
12:   {Diophantine Projection}
13:   $H_{t}\leftarrow\text{TheRing}(H_{temp})$
14:  else
15:   $H_{t}\leftarrow H_{temp}$
16:  end if
17:end for
18:Stage 4: Coda
[…]
21:return $\text{Softmax}(logits)$
Algorithm 4 Halo-N Ring Mechanism (Diophantine Solver)
1:Function TheRing($H_{\mathbb{Q}}$)
2: {Projects state onto the Cartesian Farey Lattice $\mathcal{F}^{d}$}
[…]
4: {1. Coordinate-wise Decomposition (MIMD)}
5:for all element $h_{ij}$ in tensor $H_{\mathbb{Q}}$ do
6:  $x_{val}\leftarrow h_{ij}.num/h_{ij}.denom$
7:  {2. Best Rational Approximation (Continued Fraction)}
8:  {Find $p/q$ minimizing $|x_{val}-p/q|$ s.t. $q\leq D_{max}$}
9:  $p,q\leftarrow\text{TruncatedEuclidean}(h_{ij}.num,h_{ij}.denom,D_{max})$
10:  $H_{new}[i,j]\leftarrow\text{Rational}(p,q)$
11:end for
12:return $H_{new}$
13:End Function
14:Function RationalAttention($Q,K,V$)
[…]
17:return $A\cdot V$
18:End Function
@@ 4.5 Training Objective: The Rational Alignment Loss
Unlike standard Vector Quantization (VQ) which maps vectors to an arbitrary learned codebook [Van Den Oord et al., 2017], Halo maps vectors to the fixed algebraic field of logical rationals $\mathbb{Q}_{K}=\{\frac{n}{d}\mid|d|\leq K\}$.
@@
Significance for Halo-N: While derived from the VQ-VAE paradigm used in Halo-V, this objective is equally critical for Halo-N to minimize the “Diophantine Error” during hardware resets. By strictly constraining the target space to $\mathbb{Q}_{K}$ (where denominators are bounded by $K$), $\mathcal{L}_{ring}$ acts as a complexity penalty. It actively discourages the network from generating high-entropy states that would require massive denominators to represent, effectively “pulling” potentially long, irreducible rational outputs towards lower-complexity, shorter-length rationals (e.g., simple fractions) that fit within the hardware’s optimal precision window.
4.6 Transcending the Irrational: Rational Approximations
Rational Softmax via Constructive Definitions. The standard attention mechanism requires computing $A=\text{softmax}(QK^{T}/\sqrt{d})$. We first replace the scaling factor $1/\sqrt{d}$ with an integer shift $2^{-k}$. For the exponential function, we strictly adhere to its mathematical definition as the limit of an infinite rational series. For any $x\in\mathbb{Q}$, the partial sum constitutes a dense rational approximation:
$\text{RatExp}(x,N)=\sum_{k=0}^{N}\frac{x^{k}}{k!}\xrightarrow{N\to\infty}e^{x}$ (9)
where $N$ is a dynamic precision parameter determined by the EIU. Crucially, this expansion is not merely an approximation technique but the constructive definition of the exponential function itself. Unlike floating-point implementations which rely on hardware-specific lookup tables with fixed error, RatExp allows us to trade compute time for arbitrary precision. The Rational Softmax is then defined as:
Since both numerator and denominator are in $\mathbb{Q}$, the result remains strictly rational, effectively computing the probability distribution on a progressively refined rational lattice.
$y_{t+1}=\frac{1}{2}y_{t}\left(3-\sigma^{2}y_{t}^{2}\right)$ (11)
@@ 5.4 Scale Analysis: The 600B Parameter Regime
Results in Figure 5 reveal the “Curse of Dimensionality”: larger models are less numerically stable. This creates a paradox: The larger the model, the higher the precision required.